OpenStudy (anonymous):

find an equation of variation where y varies directly as x and y=2 when x=10, find y when x= -25

OpenStudy (anonymous):

direct variation=>x1/y1=x2/y2 x1=10 y1=2 x2=-25 find y2

OpenStudy (ajprincess):

y varies directly as x means